OpenStudy (anonymous):

Write an equation for the line described. Write it in standard and point intercept form(if possible) through (1,3), m=-2 through (-5,4), m=-3/2 i need a little explanation also, if you can give it... studying for a test.

OpenStudy (anonymous):

ok so you get those in standard form by first using this formula: \[(y - y_{1}) = m (x - x_{1})\]

OpenStudy (anonymous):

so, for the first equation, this is (y - 3) = -2(x - 1)

OpenStudy (anonymous):

y - 3 = -2x + 2

OpenStudy (anonymous):

to get it into standard form (Ax + By = 0) move 2x to the other side, giving you 2x + y - 3 = 2

OpenStudy (anonymous):

depending on qhat the quiz asks you, the final answer will either be 2x + y = 5 or 2x + y - 5 = 0.

OpenStudy (anonymous):

*what

OpenStudy (anonymous):

Did you mean slope intercept when you said point intercept?

OpenStudy (anonymous):

yes.. sorry

OpenStudy (anonymous):

So the second one. Standard form : y-4=-3/2(x-(-5))

OpenStudy (anonymous):

I get confused with the fractions?

OpenStudy (anonymous):

And all those negatives are throwing me off..

OpenStudy (anonymous):

Lots of people have trouble with it. We know that a negative negative is a positive. That gives you y - 4 = -3/2(x + 5)

OpenStudy (anonymous):

Then, you multiply x and 5 by -3/2

OpenStudy (anonymous):

a positive and a negative multiplied together is a negative. That gives you -3/2x - 15/2

OpenStudy (anonymous):

-15/2 is reached when you multiply -3/2 and 5/1

OpenStudy (anonymous):

do you understand that?

OpenStudy (anonymous):

Yes, so far so good.

OpenStudy (anonymous):

Then I need to put it in standard form..

OpenStudy (anonymous):

Well, that wasn't a complete point slope form, so let me explain. Above you gave me y - 4 = -3/2(x - (-5)) Your complete slope form would be y = -3/2(x - 1), getting the 4 on the correct side. So, we'd have -3/2x + 3/2

OpenStudy (anonymous):

do you understand how I got that?

OpenStudy (anonymous):

Why dont you want it in the same form as the previous one? ax=by=0?

OpenStudy (anonymous):

sorry ax+by=0

OpenStudy (anonymous):

When I convert 4 into a fraction is that 8/2?

OpenStudy (anonymous):

When I subtract that 8/2 from the right side I come up with -3/2x-7/2?

OpenStudy (anonymous):

hold on a second

OpenStudy (anonymous):

it's easier to set up standard form if you first have it in slope intercept form.

OpenStudy (anonymous):

no, you want to get the constant on the right side of the equation BEFORE you multiply the whole thing by -3/2.

OpenStudy (anonymous):

y = -3/2 + 3/2

OpenStudy (anonymous):

constant meaning?

OpenStudy (anonymous):

Constant is a number with no variables; 3a would be a coefficient, 3 is a constant.

OpenStudy (anonymous):

where and how are you getting that +3/2

OpenStudy (anonymous):

ok, let me explain

OpenStudy (anonymous):

y-4=-3/2(x=5) can you take me from there?

OpenStudy (anonymous):

Okay. (you were right, I was getting confused.)

OpenStudy (anonymous):

(x+5)

OpenStudy (anonymous):

first you add 4 to each side, giving you y = -3/2(x + 9)

OpenStudy (anonymous):

multiply both terms by -3/2, giving you -3/2x -27/2

OpenStudy (anonymous):

move -3/2x to the left side of the equation, giving you 3/2x + y = 27/2

OpenStudy (anonymous):

sorry, -27/2

OpenStudy (anonymous):

depending on the curriculum, your final answer would be 3/2x + y = -27/2 or 3/2x + y + 27/2 = 0

OpenStudy (anonymous):

let me rewrite it out myself and see if that's the right answer

OpenStudy (anonymous):

My book gives me an answer of 3x+2y=-7

OpenStudy (anonymous):

Do I just multiply by 2 to clear the fraction?

OpenStudy (anonymous):

no, when you're talking about lines you don't take out the fraction

OpenStudy (anonymous):

but let me work it out and see if I can get it that way multiplying by 2

OpenStudy (anonymous):

ok, it seems to me that your book does want you to multiply by 2, but my curriculum didn't teach me that

OpenStudy (anonymous):

Ok.. thank you so much for being patient with me...this helped!

OpenStudy (anonymous):

you're welcome :D